# Towards Evaluating and Building Versatile Large Language Models for Medicine

### v1

@@ Abstract
In this study, we present MedS-Bench, a comprehensive benchmark designed to evaluate the performance of large language models (LLMs) in clinical contexts. Unlike existing benchmarks that focus on multiple-choice question answering, MedS-Bench spans 11 high-level clinical tasks, including clinical report summarization, treatment recommendations, diagnosis, named entity recognition, and medical concept explanation, among others. We evaluated six leading LLMs, e.g., MEDITRON, Mistral, InternLM 2, Llama 3, GPT-4, and Claude-3.5 using few-shot prompting, and found that even the most sophisticated models struggle with these complex tasks. To address these limitations, we developed MedS-Ins, a large-scale instruction tuning dataset for medicine. MedS-Ins comprises 58 medically oriented language corpora, totaling 13.5 million samples across 122 tasks. To demonstrate the dataset’s utility, we conducted a proof-of-concept experiment by performing instruction tuning on a lightweight, open-source medical language model. The resulting model, MMedIns-Llama 3, significantly outperformed existing models across nearly all clinical tasks. To promote further advancements in the application of LLMs to clinical challenges, we have made the MedS-Ins dataset fully accessible and invite the research community to contribute to its expansion. Additionally, we have launched a dynamic leaderboard for MedS-Bench, which we plan to regularly update the test set to track progress and enhance the adaptation of general LLMs to the medical domain. Leaderboard: https://henrychur.github.io/MedS-Bench/. Github: https://github.com/MAGIC-AI4Med/MedS-Ins.
@@ 1 Introduction
To advance the development of open-source medical LLMs capable of tackling a broad spectrum of clinical tasks, we take inspiration from the idea of Super-NaturalInstructions [46], and construct the first, comprehensive instruction tuning dataset for medicine, MedS-Ins. It aggregates 58 open-source biomedical natural language processing datasets from five text sources, including exams, clinical texts, academic papers, medical knowledge bases, and daily conversations, resulting in 13.5 million samples across 122 clinical tasks, each accompanied with hand-written task definitions. We performed extensive instruction tuning on open-source medical language models, and explored both zero-shot and few-shot prompting strategies. The outcome is a new medical LLM – MMedIns-Llama 3, for the first time, showing the effectiveness of training on diverse medical tasks through instruction tuning, enabling open-source medical LLMs to surpass leading closed-source models, including GPT-4 and Claude-3.5, across a wide range of clinical tasks.
@@ 2.1 The Description of MedS-Bench
To evaluate the capabilities of various LLMs in clinical applications, we developed MedS-Bench, a comprehensive medical benchmark that extends beyond traditional multiple-choice questions. MedS-Bench encompasses 11 high-level clinical task categories, derived from 39 existing datasets, as illustrated in Figure 1. Each dataset was reformatted into an instruction-prompted question-answering structure, complete with hand-crafted task definitions (instructions), as shown in Figure 2a. The task categories we considered include: Multi-choice Question Answering, Text Summarization, Information Extraction, Explanation, Rationale, Named Entity Recognition, Diagnosis, Treatment Planning, Clinical Outcome Prediction, Text Classification, Fact Verification, and Natural Language Inference. A more detailed description of each category is provided in Supplementary Sec. A.2.
@@
Figure 1: Benchmark Statistics. The hierarchical ring chart meticulously displays the data distribution within the evaluation benchmarks. The first tier categorizes the types of tasks, with the benchmarks encompassing 11 primary task categories. The second tier outlines the datasets involved, including 39 datasets in total. The third tier details the specific tasks, with the benchmarks collectively addressing 52 distinct tasks. Overall, this benchmark allows for a thorough and comprehensive evaluation of model performance across multiple dimensions.
@@ 2.3 Quantitative Results on Various Tasks
For each task type, we start by discussing the performance of various existing LLMs, followed by a comparison with our final model, MMedIns-Llama 3. All results presented here were obtained using a 3-shot prompting strategy (more details in Supplementary Sec. A.1), except for MCQA tasks, where we used a zero-shot prompting setting to align with previous studies [5, 48, 32]. As our comparisons include proprietary models like GPT-4 and Claude 3.5, which incur usage costs, we randomly sampled 50-100 test cases per benchmark to manage the cost constraints. The task description and specific sampling numbers are detailed in Supplementary Sec. A.2. For simplicity, the percentage mark (%) is omitted in all the following tables and results analysis.
@@
On these multi-choice question answering datasets, existing proprietary LLMs have demonstrated very high accuracies, for example, on MedQA, GPT-4 can achieve 85.8, which is almost comparable to human experts, and Llama 3 can also pass the exam with 60.9 scores. Similarly, in languages other than English, LLMs also demonstrate superior results in multiple-choice accuracy on MMedBenh [32]. The results indicate that as multi-choice questions have been extensively considered in existing research, different LLMs may have been specifically optimized for such tasks, resulting in high performance. It is therefore essential to build up a more comprehensive benchmark, to further push the development of LLMs towards clinical applications.
@@
[] Method MedQA MedMCQA PubMedQA MMedBench Avg. ZH JA FR RU ES Close-source Models GPT-3.5 57.7 72.7 53.8 52.3 34.6 32.5 66.4 66.1 54.5 GPT-4 85.8 72.3 70.0 75.1 72.9 56.6 83.6 85.7 75.3 Open-source Models MEDITRON 47.9 59.2 74.4 61.9 40.2 35.1 67.6 53.3 54.9 InternLM 2 - - - 77.6 47.7 41.0 68.4 59.6 - Mistral 50.8 48.2 75.4 71.1 44.7 48.7 74.2 63.9 49.1 Llama 3 60.9 50.7 73.0 78.2 48.2 50.8 71.5 64.2 62.2 MMedIns-Llama 3 63.6 57.1 78.2 78.6 54.3 46.0 72.3 61.2 63.9
Table 1: Results on medical multiple-choice question answering, as reported with Accuracy score.
@@
Table 2: Results on text summarization, as reported with BLEU and ROUGE scores, formatted as ‘BLEU/ROUGE’. “Con.” denotes conclusions.
@@
The performance of information extraction is summarized in Table 3. InternLM 2 shows exceptionally good performance in this task with an average score of 81.58. For example, In the PICO tasks, InternLM 2 leads in both Intervention and Outcome Extraction, with scores of 79.07 and 76.74, respectively. Closed-source models such as GPT-4 and Claude-3.5 outperform all other open-source counterparts, with average scores of 77.49 and 78.86, respectively. Analysis of individual benchmark components reveals that most LLMs perform better at extracting less medically complex information, such as basic patient details, compared to more specialized medical data like outcomes and interventions. For instance, in extracting basic information from PMC Patients, most LLMs score above 90, with Claude-3.5 achieving the highest score of 98.02. In contrast, performance on Clinical Outcome Extraction tasks within PICO is relatively poor.
Our proposed model, MMedIns-Llama 3, demonstrates the best overall performance, achieving an average score of 83.18, surpassing InternLM 2 by 1.6 points. Notably, in the PICO tasks, MMedIns-Llama 3 excels in Participant Extraction, scoring 83.72, which exceeds the second-best model by 9.3 points.
[] Method PICO ADE PMC patient Avg. Participant Ext. Intervention Ext. Outcome Ext. Drug Dose Ext. Basic Info. Ext. Close-source Models GPT-4 72.73 72.09 67.44 79.17 96 77.49 Claude-3.5 67.44 74.42 67.44 86.96 98 78.86 Open-source Models MEDITRON 62.79 53.49 60.47 73.91 68 63.73 InternLM 2 74.42 79.07 76.74 95.65 82 81.58 Llama 3 65.12 74.42 60.47 91.3 94 77.06 Mistral 62.79 62.79 58.14 91.3 90 73.01 MMedIns-Llama 3 83.72 76.74 62.79 95.65 97 83.18
Table 3: Results on information extraction, as reported with Accuracy score. “Ext.” denotes extraction and “Info.” denotes information.
@@
In Table 4, we evaluate the model on medical concept explanation, GPT-4, Llama 3, and Mistral perform well on this task, achieving an average of 19.18/21.33, 21.14/23.33, and 19.18/21.33 respectively. On the contrary, the scores of Claude-3.5, InternLM 2, and MEDITRON are relatively lower, with averages of 14.78/19.29, 13.56/18.93, and 11.21/22.66, respectively. We conjecture that the relatively poor performance of MEDITRON can be attributed to its training corpus, which is more focused on academic papers and guidelines, and thus falls short in basic medical concept explanation.
Our final model, MMedIns-Llama 3, significantly outperforms the other ones across all concept explanation tasks, particularly in Health Fact Explanation (35.56/32.32) and BioLORD Concept Explanation (32.36/40.39), and achieving the highest average scores of 33.89/37.68. Following MMedIns-Llama 3, Llama 3 and GPT-4 also showed strong performance, with Llama 3 scoring 21.14/23.33 and GPT-4 closely behind at 19.18/21.33.
[] Method Health Fact Exp. Do Entity Exp. BioLORD Concept Exp. Avg. Close-source Models GPT-4 18.27/20.45 20.69/22.85 18.59/20.69 19.18/21.33 Claude-3.5 14.55/19.61 11.74/17.08 18.05/21.19 14.78/19.29 Open-source Models MEDITRON 12.48/18.83 8.17/25.89 12.96/23.27 11.21/22.66 InternLM 2 22.40/27.18 6.28/12.69 12.01/16.91 13.56/18.93 Llama 3 20.47/21.55 20.94/24.13 22.01/24.33 21.14/23.33 Mistral 17.81/24.00 17.34/21.66 22.18/22.79 19.11/22.82 MMedIns-Llama 3 35.56/32.32 33.74/40.34 32.36/40.39 33.89/37.68
Table 4: Results on medical concept explanation, as reported with ‘BLEU/ROUGE’ scores. “Exp.” denotes Explanation
@@
In Table 5, we evaluate the complex rationale, i.e., explaining the answer, and comparing the reasoning abilities of various models using the MMedBench [32] dataset across six languages. Among the models tested, the closed-source model Claude-3.5 exhibited the strongest performance, with average scores of 46.03/37.65, demonstrating consistently high scores across all languages, particularly in French and Spanish. This superior performance may be attributed to the similarity of this task to chain-of-thought reasoning, a capability that has been specifically enhanced in many general-purpose LLMs. Among open-source models, Mistral and InternLM 2 showed comparable performance, with average scores of 37.61/31.55 and 30.03/26.44, respectively. It is important to note that GPT-4 was excluded from this evaluation, because the rationale component of the MMedBench dataset was primarily constructed using GPT-4 outputs, which could introduce bias and bring unfair comparisons.
Consistent with our observations in concept explanation, our final model, MMedIns-Llama 3, demonstrated the best overall performance with average BLEU-1/ROUGE-1 scores of 47.17/34.96 across all languages, notably, achieving 53.31/36.83 in Japanese reasoning tasks, 48.54/38.22 in English, and 47.64/39.24 in French, respectively. This superior performance is likely due to the fact that our base language model (MMed-Llama 3) was initially developed to be multilingual [32]. Consequently, even though our instruction tuning did not explicitly target multilingual data, the final model outperforms others across multiple languages.
[] Method MMedBench Avg. Chinese English French Japanese Russian Spanish Close-source Models Claude-3.5 45.77/35.79 46.37/39.57 49.28/41.04 46.65/40.56 38.49/28.50 49.64/40.45 46.03/37.65 Open-source Models MEDITRON 20.18/20.76 39.50/30.88 35.01/29.36 21.41/26.46 26.56/20.99 37.49/30.17 30.03/26.44 InternLM 2 36.66/30.52 44.62/39.36 36.50/34.21 29.10/34.77 27.97/23.24 41.97/36.13 36.14/33.04 Llama 3 33.71/27.83 44.86/38.75 23.68/22.22 9.44/12.92 31.92/24.69 32.33/28.04 29.32/25.74 Mistral 34.98/28.34 47.29/37.06 38.48/35.04 29.08/27.50 32.10/24.94 43.74/36.39 37.61/31.55 MMedIns-Llama 3 51.50/35.35 48.54/38.22 47.64/39.24 53.31/36.83 34.91/23.71 47.08/36.42 47.17/34.96
Table 5: Results on rationale, as reported with ‘BLEU/ROUGE’ scores. Note here, we do not include the results for GPT-4 since the original evaluation sets were generated with it, that may bring unfair comparison bias.
@@
As shown in Table 6, among the six existing models, GPT-4 is the only one that consistently demonstrates robust performance across Named Entity Recognition (NER) tasks, achieving an average F1-Score of 44.30. It excels particularly in the BC5Chem Chemical Recognition task with a scores of 63.77. InternLM 2 also delivers competitive results, with an average F1-Score of 40.81, showing strong performance in both the BC5Chem and BC5Disease tasks. Llama 3 and Mistral, with average F1-Scores of 24.70 and 20.10, respectively, exhibit moderate performance. MEDITRON, not optimized for NER tasks, shows limited effectiveness in this area.
Our model, MMedIns-Llama 3, significantly outperforms all other models, achieving an an average F1-Score of 68.58. It excels in the BC4Chem and BC5Chem Chemical Recognition tasks, with F1 scores of 78.62 and 89.71, respectively. Furthermore, MMedIns-Llama 3 leads in the BC5Disease Disease Recognition task with an F1-Score of 51.97 and in the Species800 Organism Recognition task with 54.02, demonstrating superior capability in handling complex NER tasks across various biomedical domains.
Table 6: Results on NER tasks, as reported with F1-Score scores. ‘Rec.’ is short for ‘recognition’.
     Method      BC4Chem      BC5Chem      BC5Disease      Species800      Avg.
     Chemical Rec.       Disease Rec.      Organism Rec.
     Close-source Models
     GPT-4      47.41      63.77      35.77      30.26      44.30
     Claude-3.5      30.77      48.61      27.14      9.88      29.10
     Open-source Models
     MEDITRON      0.83      6.53      2.45      0.80      2.65
     InternLM 2      32.79      50.39      42.02      38.03      40.81
     Llama 3      24.62      42.55      24.00      7.64      24.70
     Mistral      21.48      41.48      13.25      4.21      20.10
     MMMedIns-Llama 3      78.62      89.71      51.97      54.02      68.58
@@
Overall, the open-source LLMs underperform the closed-source counterparts in these tasks, and in some instances, they fail to provide meaningful predictions. For example, Llama 3 struggles with predicting critical triage and the 72-hour ED revisit binary indicator. In the SEER treatment planning task, InternLM 2 and MEDITRON achieve moderate accuracy scores of 68 and 60, respectively. For the DDXPlus diagnosis task, InternLM 2 and Llama 3 perform slightly better, with an accuracy of 32. However, closed-source models such as GPT-4 and Claude-3.5 demonstrate significantly better performance. Claude-3.5, for instance, achieves a 90 accuracy score in treatment planning, while GPT-4 attains 52 in diagnosis, highlighting the considerable gap between open-source and closed-source LLMs. Despite these results, the scores remain insufficient for reliable clinical use.
In contrast, MMedIns-Llama 3 demonstrate superior accuracy in clinical decision support tasks, with a 98 accuracy score in treatment planning, 95 in diagnosis, and an average accuracy of 86.67 (mean on the scores of Hospitalization, 72h ED Revisit, and Critical Triage) in clinical outcome prediction.
[] Method SEER DDXPlus MIMIC4ED HoC Classification Hospitalization 72h ED Revisit Critical Triage Precision Recall F1 Close-source Models GPT-4 78 52 62 69 66 55.07 72.37 60.38 Claude-3.5 90 48 64 63 57 55.18 80.96 63.32 Open-source Models MEDITRON 60 24 36 4 11 23.39 36.85 26.21 InternLM 2 68 32 60 76 69 22.97 83.16 32.72 Llama 3 52 32 51 4 3 26.6 69.59 36.18 Mistral 36 28 58 39 43 35.18 59.27 40.80 MMedIns-Llama 3 98 95 69 97 94 91.29 85.57 87.37
Table 7: Results on treatment planning, diagnosis clinical outcome prediction, and text classification results. The first 3 tasks are reported with Accuracy score, and text classification is reported with Precision, Recall, and F1 score.
@@
GPT-4 and Claude-3.5 perform well on this task, with GPT-4 achieving a macro-F1 score of 60.38 and Claude-3.5 slightly better at 63.32. Both models show strong recall capabilities, particularly Claude-3.5, which achieves a macro-Recall of 80.96, underscoring its proficiency in identifying relevant labels. Mistral shows moderate performance with a macro-F1 score of 40.8, maintaining a balance between precision and recall. In contrast, Llama 3 and InternLM 2 display lower overall performance, with macro-F1 scores of 36.18 and 32.72, respectively. These models, particularly InternLM 2, demonstrate high recall but struggle with precision, resulting in lower F1 scores. MEDITRON ranks lowest in this task, with a macro-F1 score of 26.21.
MMedIns-Llama 3 clearly outperforms all other models, achieving the highest scores across all metrics, with a macro-Precision of 91.29, a macro-Recall of 85.57, and a macro-F1 score of 87.37. These results highlight MMedIns-Llama 3’s superior ability to accurately classify and recall multiple labels, making it the most effective model for this complex task.

### v2

@@ Abstract
In this study, we present MedS-Bench, a comprehensive benchmark designed to evaluate the performance of large language models (LLMs) in clinical contexts. Unlike existing benchmarks that focus on multiple-choice question answering, MedS-Bench spans 11 high-level clinical tasks, including clinical report summarization, treatment recommendations, diagnosis, named entity recognition, and medical concept explanation, among others. We evaluated six leading LLMs, e.g., MEDITRON, Mistral, InternLM 2, Llama 3, GPT-4, and Claude-3.5 using few-shot prompting, and found that even the most sophisticated models struggle with these complex tasks. To address these limitations, we developed MedS-Ins, a large-scale instruction tuning dataset for medicine. MedS-Ins comprises 58 medically oriented language corpora, totaling 5M instances with 19K instructions, across 122 tasks. To demonstrate the dataset’s utility, we conducted a proof-of-concept experiment by performing instruction tuning on a lightweight, open-source medical language model. The resulting model, MMedIns-Llama 3, significantly outperformed existing models across nearly all clinical tasks. To promote further advancements in the application of LLMs to clinical challenges, we have made the MedS-Ins dataset fully accessible and invite the research community to contribute to its expansion. Additionally, we have launched a dynamic leaderboard for MedS-Bench, which we plan to regularly update the test set to track progress and enhance the adaptation of general LLMs to the medical domain. Leaderboard: https://henrychur.github.io/MedS-Bench/. Github: https://github.com/MAGIC-AI4Med/MedS-Ins.
@@ 1 Introduction
To advance the development of open-source medical LLMs capable of tackling a broad spectrum of clinical tasks, we take inspiration from the idea of Super-NaturalInstructions [46], and construct the first, comprehensive instruction tuning dataset for medicine, MedS-Ins. It aggregates 58 open-source biomedical natural language processing datasets from five text sources, including exams, clinical texts, academic papers, medical knowledge bases, and daily conversations, resulting in 5M instances with 19K instructions across 122 clinical tasks, each accompanied with hand-written task definitions. We performed extensive instruction tuning on open-source medical language models, and explored both zero-shot and few-shot prompting strategies. The outcome is a new medical LLM – MMedIns-Llama 3, for the first time, showing the effectiveness of training on diverse medical tasks through instruction tuning, enabling open-source medical LLMs to surpass leading closed-source models, including GPT-4 and Claude-3.5, across a wide range of clinical tasks.
@@ 2.1 The Description of MedS-Bench
To evaluate the capabilities of various LLMs in clinical applications, we developed MedS-Bench, a comprehensive medical benchmark that extends beyond traditional multiple-choice questions. MedS-Bench encompasses 11 high-level clinical task categories, derived from 28 existing datasets, as illustrated in Figure 1. Each dataset was reformatted into an instruction-prompted question-answering structure, complete with hand-crafted task definitions (instructions), as shown in Figure 2a. The task categories we considered include: Multi-choice Question Answering, Text Summarization, Information Extraction, Explanation, Rationale, Named Entity Recognition, Diagnosis, Treatment Planning, Clinical Outcome Prediction, Text Classification, Fact Verification, and Natural Language Inference. A more detailed description of each category is provided in Supplementary Sec. A.2.
@@
Figure 1: Benchmark Statistics. The hierarchical ring chart meticulously displays the data distribution within the evaluation benchmarks. The first tier categorizes the types of tasks, with the benchmarks encompassing 11 primary task categories. The second tier outlines the datasets involved, including 28 datasets in total. The third tier details the specific tasks, with the benchmarks collectively addressing 52 distinct tasks. Overall, this benchmark allows for a thorough and comprehensive evaluation of model performance across multiple dimensions.
@@ 2.3 Quantitative Results on Various Tasks
For each task type, we start by discussing the performance of various existing LLMs, followed by a comparison with our final model, MMedIns-Llama 3. All results presented here were obtained using a 3-shot prompting strategy (more details in Supplementary Sec. A.1), except for MCQA tasks, where we used a zero-shot prompting setting to align with previous studies [5, 48, 32]. As our comparisons include proprietary models like GPT-4 and Claude 3.5, which incur usage costs, we randomly sampled around 1500 test cases per benchmark to manage the cost constraints. The task description and specific sampling numbers are detailed in Supplementary Sec. A.2. For simplicity, the percentage mark (%) is omitted in all the following tables and results analysis.
@@
On these multi-choice question answering datasets, existing proprietary LLMs have demonstrated very high accuracies, for example, on MedQA, GPT-4 can achieve 85.8, which is almost comparable to human experts, and Llama 3 can also pass the exam with 60.9 scores. Similarly, in languages other than English, LLMs also demonstrate superior results in multiple-choice accuracy on MMedBench [32]. The results indicate that as multi-choice questions have been extensively considered in existing research, different LLMs may have been specifically optimized for such tasks, resulting in high performance. It is therefore essential to build up a more comprehensive benchmark, to further push the development of LLMs towards clinical applications.
@@
[] Method MedQA MedMCQA PubMedQA MMedBench Avg. ZH JA FR RU ES Close-source Models GPT-3.5 57.7 72.7 53.8 52.3 34.6 32.5 66.4 66.1 54.5 GPT-4 85.8 72.3 75.2 75.1 72.9 56.6 83.6 85.7 75.3 Open-source Models MEDITRON 47.9 59.2 74.4 61.9 40.2 35.1 67.6 53.3 54.9 InternLM 2 - - - 77.6 47.7 41.0 68.4 59.6 - Mistral 50.8 48.2 75.4 71.1 44.7 48.7 74.2 63.9 49.1 Llama 3 60.9 50.7 73.0 78.2 48.2 50.8 71.5 64.2 62.2 MMedIns-Llama 3 63.6 57.1 78.2 78.6 54.3 46.0 72.3 61.2 63.9
Table 1: Results on medical multiple-choice question answering, as reported with Accuracy score. Bolding represents the best results.
@@
Table 2: Results on text summarization, as reported with BLEU and ROUGE scores, formatted as ‘BLEU/ROUGE’. “Con.” denotes conclusions. Bolding represents the best results.
@@
The performance of information extraction is summarized in Table 3. InternLM 2 shows exceptionally good performance in this task with an average score of 79.11. For example, In the PICO tasks, InternLM 2 leads in both Intervention and Outcome Extraction, with scores of 74.42 and 69.77, respectively. Closed-source models such as GPT-4 and Claude-3.5 outperform all other open-source counterparts, with average scores of 76.92 and 79.41, respectively. Analysis of individual benchmark components reveals that most LLMs perform better at extracting less medically complex information, such as basic patient details, compared to more specialized medical data like outcomes and interventions. For instance, in extracting basic information from PMC Patients, most LLMs score above 90, with Claude-3.5 achieving the highest score of 99.07. In contrast, performance on Clinical Outcome Extraction tasks within PICO is relatively poor.
Our proposed model, MMedIns-Llama 3, demonstrates the best overall performance, achieving an average score of 83.77, surpassing InternLM 2 by 4.66 points. Notably, in the PICO tasks, MMedIns-Llama 3 excels in Participant Extraction, scoring 83.72, which exceeds the second-best model by 11.63 points.
[] Method PICO ADE PMC patient Avg. Participant Ext. Intervention Ext. Outcome Ext. Drug Dose Ext. Basic Info. Ext. Close-source Models GPT-4 67.44 62.79 65.12 91.30 97.93 76.92 Claude-3.5 65.12 76.74 60.47 95.65 99.07 79.41 Open-source Models MEDITRON 72.09 46.51 51.16 95.65 72.20 67.52 InternLM 2 72.09 74.42 69.77 95.65 83.60 79.11 Llama 3 58.14 79.07 58.14 69.57 95.93 72.17 Mistral 60.47 65.12 48.84 91.30 85.20 70.18 MMedIns-Llama 3 83.72 79.07 62.79 95.65 97.60 83.77
Table 3: Results on information extraction, as reported with Accuracy score. “Ext.” denotes extraction and “Info.” denotes information. Bolding represents the best results.
@@
In Table 4, we evaluate the model on medical concept explanation, GPT-4, Llama 3, and Mistral perform well on this task, achieving an average of 19.37/21.58, 13.51/17.92, and 13.53/17.37 respectively. On the contrary, the scores of Claude-3.5, InternLM 2, and MEDITRON are relatively lower, with averages of 12.56/16.75, 11.53/17.01, and 8.51/18.90, respectively. We conjecture that the relatively poor performance of MEDITRON can be attributed to its training corpus, which is more focused on academic papers and guidelines, and thus falls short in basic medical concept explanation.
Our final model, MMedIns-Llama 3, significantly outperforms the other ones across all concept explanation tasks, particularly in Health Fact Explanation (30.50/28.53) and BioLORD Concept Explanation (38.12/43.90), and achieving the highest average scores of 34.43/37.47. Following MMedIns-Llama 3, GPT-4 also showed strong performance, with GPT-4 scoring 19.37/21.58.
[] Method Health Fact Exp. Do Entity Exp. BioLORD Concept Exp. Avg. Close-source Models GPT-4 18.63/20.80 19.14/21.14 20.33/22.80 19.37/21.58 Claude-3.5 14.96/18.48 8.75/13.28 13.95/18.49 12.56/16.75 Open-source Models MEDITRON 6.09/8.65 7.68/25.39 11.76/22.66 8.51/18.90 InternLM 2 22.36/27.01 5.28/10.39 6.95/13.62 11.53/17.01 Llama 3 16.79/20.32 14.88/18.84 8.87/14.61 13.51/17.92 Mistral 18.11/21.31 9.21/14.11 13.27/16.68 13.53/17.37 MMedIns-Llama 3 30.50/28.53 34.66/39.99 38.12/43.90 34.43/37.47
Table 4: Results on medical concept explanation, as reported with ‘BLEU/ROUGE’ scores. “Exp.” denotes Explanation. Bolding represents the best results.
@@
In Table 5, we evaluate the complex rationale, i.e., explaining the answer and comparing the reasoning abilities of various models using the MMedBench [32] dataset across six languages. Among the models tested, the closed-source model Claude-3.5 exhibited the strongest performance, with average scores of 46.26/36.97, demonstrating consistently high scores across all languages, particularly in French and Spanish. This superior performance may be attributed to the similarity of this task to chain-of-thought reasoning, a capability that has been specifically enhanced in many general-purpose LLMs. Among open-source models, Mistral and InternLM 2 showed comparable performance, with average scores of 38.14/32.28 and 35.65/32.04, respectively. It is important to note that GPT-4 was excluded from this evaluation, because the rationale component of the MMedBench dataset was primarily constructed using GPT-4 outputs, which could introduce bias and bring unfair comparisons.
Consistent with our observations in concept explanation, our final model, MMedIns-Llama 3, demonstrated the best overall performance with average BLEU-1/ROUGE-1 scores of 46.90/34.54 across all languages, notably, achieving 51.74/35.19 in Japanese reasoning tasks, 49.08/38.19 in English, and 46.93/38.73 in French, respectively. This superior performance is likely due to the fact that our base language model (MMed-Llama 3) was initially developed to be multilingual [32]. Consequently, even though our instruction tuning did not explicitly target multilingual data, the final model outperforms others across multiple languages.
[] Method MMedBench Avg. Chinese English French Japanese Russian Spanish Close-source Models Claude-3.5 44.64/34.63 47.07/38.67 48.93/41.23 49.22/39.15 38.90/28.17 48.80/39.99 46.26/36.97 Open-source Models MEDITRON 20.39/21.79 38.42/31.24 34.43/29.33 18.89/24.98 24.32/16.77 37.64/31.01 29.01/25.86 InternLM 2 35.23/30.77 44.12/37.39 36.10/33.65 29.13/33.15 27.43/20.99 41.87/36.30 35.65/32.04 Llama 3 28.51/23.30 44.10/39.26 24.92/22.24 13.46/15.04 31.16/22.85 32.37/27.70 29.09/25.06 Mistral 35.53/28.91 47.20/37.88 39.53/35.64 29.16/28.96 32.15/23.99 45.27/38.33 38.14/32.28 MMedIns-Llama 3 50.27/34.01 49.08/38.19 46.93/38.73 51.74/35.19 35.27/23.81 48.15/37.35 46.90/34.54
Table 5: Results on rationale, as reported with ‘BLEU/ROUGE’ scores. Note here, we do not include the results for GPT-4 since the original evaluation sets were generated with it, that may bring unfair comparison bias. Bolding represents the best results.
@@
As shown in Table 6, among the six existing models, GPT-4 is the only one that consistently demonstrates robust performance across Named Entity Recognition (NER) tasks, achieving an average F1-Score of 59.52. It excels particularly in the BC5Chem Chemical Recognition task with a score of 67.62. InternLM 2 also delivers competitive results, with an average F1-Score of 45.69, showing strong performance in both the BC5Chem and BC5Disease tasks. Llama 3 and Mistral, with average F1-Scores of 23.62 and 16.53, respectively, exhibit moderate performance. MEDITRON, not optimized for NER tasks, shows limited effectiveness in this area.
Our model, MMedIns-Llama 3, significantly outperforms all other models, achieving an an average F1-Score of 79.29. It excels in the BC4Chem and BC5Chem Chemical Recognition tasks, with F1 scores of 90.78 and 91.25, respectively. Furthermore, MMedIns-Llama 3 leads in the BC5Disease Disease Recognition task with an F1-Score of 54.26 and in the Species800 Organism Recognition task with 80.87, demonstrating superior capability in handling complex NER tasks across various biomedical domains.
[]      Method      BC4Chem      BC5Chem      BC5Disease      Species800      Avg.      Chemical Rec.       Disease Rec.      Organism Rec.      Close-source Models      GPT-4      54.84      67.62      53.20      62.43      59.52      Claude-3.5      22.98      40.77      24.05      14.45      25.56      Open-source Models      MEDITRON      1.98      4.11      1.33      0.40      1.96      InternLM 2      41.21      41.51      37.11      62.93      45.69      Llama 3      19.45      37.83      25.30      11.90      23.62      Mistral      15.56      32.09      12.17      6.31      16.53      MMedIns-Llama 3      90.78      91.25      54.26      80.87      79.29
Table 6: Results on NER tasks, as reported with F1-Score scores. ‘Rec.’ is short for ‘recognition’. Bolding represents the best results.
@@
Overall, the open-source LLMs underperform the closed-source counterparts in these tasks, and in some instances, they fail to provide meaningful predictions. For example, Llama 3 struggles with predicting critical triage and the 72-hour ED revisit binary indicator. In the SEER treatment planning task, InternLM 2 and MEDITRON achieved moderate accuracy scores of 62.33 and 68.27, respectively. For the DDXPlus diagnosis task, InternLM 2 and Mistral perform slightly better, with an accuracy of 35.20 and 34.80 respectively. Notably, for clinical outcome prediction, we find Llama 3 fails in this task. It appears that because this task is outside its training distribution, it often fails to follow the three-shot provided format, thus resulting in extremely low scores, 9.27 and 8.80 in predicting 72h ED Revisit and Critical Triage respectively. However, closed-source models such as GPT-4 and Claude-3.5 demonstrate significantly better performance. Claude-3.5, for instance, achieves a 92.93 accuracy score in treatment planning and GPT-4 attains 84.73. They also demonstrate better performance in diagnosis, highlighting the considerable gap between open-source and closed-source LLMs. Despite these results, the scores remain insufficient for reliable clinical use.
In contrast, MMedIns-Llama 3 demonstrate superior accuracy in clinical decision support tasks, with a 98.47 accuracy score in treatment planning, 97.53 in diagnosis, and an average accuracy of 63.35 (mean on the scores of Hospitalization, 72h ED Revisit, and Critical Triage) in clinical outcome prediction.
[] Method SEER DDXPlus MIMIC4ED HoC Classification Hospitalization 72h ED Revisit Critical Triage Precision Recall F1 Close-source Models GPT-4 84.73 58.13 61.20 58.07 60.13 61.07 80.23 68.06 Claude-3.5 92.93 60.24 65.80 57.91 68.53 58.43 79.84 66.74 Open-source Models MEDITRON 68.27 29.53 56.27 48.47 45.67 19.61 34.61 23.70 InternLM 2 62.33 35.20 58.80 55.13 52.80 20.65 82.24 31.09 Llama 3 56.07 33.73 39.07 9.27 8.80 32.40 52.03 38.37 Mistral 38.93 34.80 56.27 48.47 45.67 40.39 64.11 48.73 MMedIns-Llama 3 98.47 97.53 74.20 52.73 63.13 89.59 85.58 86.66
Table 7: Results on treatment planning, diagnosis clinical outcome prediction, and text classification results. The first 3 tasks are reported with Accuracy score, and text classification is reported with Precision, Recall, and F1 scores. Bolding represents the best results.
@@
GPT-4 and Claude-3.5 perform well on this task, with GPT-4 achieving a macro-F1 score of 68.06 and Claude-3.5 slightly worse at 66.74. Both models show strong recall capabilities, particularly GPT-4, which achieves a macro-Recall of 80.23, underscoring its proficiency in identifying relevant labels. Mistral shows moderate performance with a macro-F1 score of 48.73, maintaining a balance between precision and recall. In contrast, Llama 3 and InternLM 2 display lower overall performance, with macro-F1 scores of 38.37 and 31.09, respectively. These models, particularly InternLM 2, demonstrate high recall but struggle with precision, resulting in lower F1 scores. MEDITRON ranks lowest in this task, with a macro-F1 score of 23.70.
MMedIns-Llama 3 clearly outperforms all other models, achieving the highest scores across all metrics, with a macro-Precision of 89.59, a macro-Recall of 85.58, and a macro-F1 score of 86.66. These results highlight MMedIns-Llama 3’s superior ability to accurately classify and recall multiple labels, making it the most effective model for this complex task.